OpenStudy (anonymous):

The bearcat school store is running a promotion on school supplies .Difference items are placed on two shelves .You can purchase 3 items from shelf M and 2 from shelf N for $16.Or you can purchase 2 item from shelf M and 3 from shelf M for $14 .If the supplies on shelf M are normally $ 6 each and the supplies on shelf N are normally $ 3, how much will you save on each package pan from part M ? 3M+2N = 16 2M+3N = 14

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

that the equation , but how I answer the question?

OpenStudy (anonymous):

not finsh yet

OpenStudy (anonymous):

first plug in the N, so you can find the M

OpenStudy (anonymous):

give M=6 N = 3

OpenStudy (anonymous):

then subtract M from the original price of M to find the save you are making

OpenStudy (anonymous):

2M + 3(3) = 14 2M + 9 = 14 2M = 5 M = 2.5 (Original M) - 2.5 = saving

OpenStudy (anonymous):

I did the second equation

OpenStudy (anonymous):

Now do the first.

OpenStudy (anonymous):

Original M = 6

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

but the given give both M and N M=6 N = 3 you don't plus M?

OpenStudy (anonymous):

Wait, sorry

OpenStudy (anonymous):

I gave you wrong info

OpenStudy (anonymous):

Why not do the elimination method to find M

OpenStudy (anonymous):

to both equations

OpenStudy (anonymous):

if you do elimination method. M =4

OpenStudy (anonymous):

\[3M+2N = 16 \\] \[2M+3N = 14\]

OpenStudy (anonymous):

the resut I did M=4 N = 2 but the given give both M and N M=6 N = 3 I think the question don't want use the resut, requite use M=6 , N=3

OpenStudy (anonymous):

N, the question wants to know what the saving are from M on the package deals.

OpenStudy (anonymous):

and you found that M = 4 on the package deals

OpenStudy (anonymous):

So, 2 dollars would be the price you saved for M

OpenStudy (anonymous):

yes I did get answer : M=4 N = 2 but the question gien: M=6 N=3

OpenStudy (anonymous):

what you do with : M=6 N =3 ?